Consider a file system that uses inodes to represent files. Diskblocks are 8 kb in size, and a pointer to a disk block requires 4bytes. this file system has 12 direct disk blocks, as well assingle, double, and triple in direct disk blocks. What is themaximum size of a file that can be stored in this file system?

Explain a little bit^^ Thank you
Answer #1

Size of the disk block \(=8 \mathrm{~KB}\)

$$ \begin{aligned} &=8 \times 1024 \text { bytes } \\ &=8192 \text { bytes } \end{aligned} $$

Number of bytes required for a pointer to point to a disk block \(=4\) bytes

Number of pointers per block \(=\frac{8 \mathrm{~KB}}{4}\)

$$ \begin{aligned} &=\frac{8192}{4} \\ &=2048 \end{aligned} $$

Number of direct disk blocks \(=12\)

As the file has 12 direct disk blocks, plus single, double, and triple indirect

disk blocks,

$$ \begin{aligned} \text { Maximum size } &=(12 \times 8 / \mathrm{KB})+(2048 \times 8 / \mathrm{KB})+(2048 \times 2048 \times 8 / \mathrm{KB}) \\ &+(2048 \times 2048 \times 2048 \times 8 / \mathrm{KB}) \\ &=68753047648 \mathrm{~KB} \end{aligned} $$

\(=64\) terabytes

Hence, the maximum size of a file that can be stored in is file system is 64 terabytes.